OpenStudy (anonymous):

Hi! Good afternoon. I'm Maynard E. Javier student in SLSU. I have a project. Could anyone help me? Can you explain to me the Area of Polar Coordinates?

OpenStudy (anonymous):

Are you @khel ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Oh, sorry. khel asked me to help him/her with project, I assumed you are the same person. Besides you can google it. https://www.google.com/search?q=Area+of+Polar+Coordinates&aq=f&oq=Area+of+Polar+Coordinates&aqs=chrome.0.57&sourceid=chrome&ie=UTF-8

OpenStudy (anonymous):

No, I'm not. But thank you anyway.

OpenStudy (sirm3d):

anong kurso mo maynard?

OpenStudy (anonymous):

I am solving area of polar coordinates and I encounter a problem. I cannot understand why (1-cos2θ)dθ becomes [θ-1/2sin2θ] from the equation 〖(1+sinθ)〗^2 dθ. Can you please answer me. Thank you.

OpenStudy (sirm3d):

@dranyam , anong kurso mo?

OpenStudy (anonymous):

BS Math po

OpenStudy (sirm3d):

ah, kaklase mo ung pauline par?

OpenStudy (anonymous):

yes. why? do we have the same topic?

OpenStudy (sirm3d):

di po. points in space topic nya. as to your question, \[\int (1-\cos2\theta) d\theta=\int 1 d\theta - \int \cos 2\theta d\theta\]

OpenStudy (sirm3d):

standard formula \[\Large \int \cos ax dx = \frac{1}{a}\sin ax + C\]

OpenStudy (sirm3d):

nasagot na po ba?

OpenStudy (anonymous):

yes. thank you very much. How about the intersection of the graph? I saw in a certain example that the interval of the equation r=2-sinΘ and r=3sinΘ is from Π/6 to Π/2. How should I know the interval? Is there any other formula?

OpenStudy (sirm3d):

can you be more specific with what is asked involving two polar equations?

OpenStudy (anonymous):

How can I get the interval of that two equations?

OpenStudy (sirm3d):

solve the system of equations \[\large \begin{cases}r=2-\sin\theta\\r=3\sin\theta\end{cases}\] here's the graph courtesy of Microsoft Mathematics the two points of intersection are at \[(\frac{3}{2},\frac{\pi}{6}),\quad(\frac{3}{2},\frac{5\pi}{6})\], not at pi/2